OpenStudy (anonymous):

there are 4 teams in a school tournament. Team A has a 25% chance of winning. Team B has the same chance as Team D. Team C has half the chance of winning as Team B. Create a table of probabilities for the sample space.

OpenStudy (anonymous):

So we know that A=.25, B=D, and C=.5B. We also know that A + B + C + D = 1, because someone has to win. A + B + C + D = 1 can be written as A + B + .5B + B = 1 .25 + 2.5B = 1 B = .75/2.5 B = .3 Using value of B you get A=.25, B=.3, C=.15 and D=.3 .25 + .3 + 1.5 + .3 = 1

OpenStudy (anonymous):

thank you Pete but can you write this in fractions?

OpenStudy (anonymous):

Sure! .25 is 1/4, .3 is 3/10, and .15 is 3/20. :) A=1/4, B=3/10, C=3/20, D=3/10

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Not a problem. :) Have a good one!

OpenStudy (anonymous):

you too!